OpenStudy (anonymous):

a bag contains 2 green,9 brown, 7 yellow, and 4 blue marbles help with a few of the probabilitys?

OpenStudy (anonymous):

p(green then blue) p(blue then blue) p(brown then not green)???? and once a marble is selected it is not replaced thank u so much

OpenStudy (anonymous):

what did you do?

OpenStudy (anonymous):

im confused with how to get it for more thn 1

OpenStudy (amoodarya):



OpenStudy (anonymous):

P(green then blue). what is P(green)?

OpenStudy (anonymous):

permutation or combination that means

OpenStudy (anonymous):

permutation. Green then blue is different from blue then green

OpenStudy (anonymous):

woulnt it be 6/22 ???

OpenStudy (anonymous):

correct. and because you can not replace the ball that is picked out, how many are left?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

shouldn't there be 21 balls left? you had 22 initially

OpenStudy (anonymous):

so it would be out of 21?

OpenStudy (anonymous):

wait, why 6/22? what's that?

OpenStudy (anonymous):

green+blue/ all the marbles

OpenStudy (anonymous):

we're looking for P(green), which is 2/22, am correct?

OpenStudy (anonymous):

green, then blue yes

OpenStudy (anonymous):

P(blue) = 4/21

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

yes we r finding it foe green, then blue the probability of picking them

OpenStudy (anonymous):

so P(green then blue) = 2/22 * 4/21 <--

OpenStudy (anonymous):

so 5.25 ah thank u

OpenStudy (anonymous):

you did the math wrong. (2/22) * (4/21) = 0.01731601731

OpenStudy (anonymous):

thx u sm